# CoreEditor: Consistent 3D Editing via Correspondence-constrained Diffusion

### v1

@@ Abstract
Text-driven 3D editing seeks to modify 3D scenes according to textual descriptions, and most existing approaches tackle this by adapting pre-trained 2D image editors to multi-view inputs. However, without explicit control over multi-view information exchange, they often fail to maintain cross-view consistency, leading to insufficient edits and blurry details. We introduce CoreEditor, a novel framework for consistent text-to-3D editing. The key innovation is a correspondence-constrained attention mechanism that enforces precise interactions between pixels expected to remain consistent throughout the diffusion denoising process. Beyond relying solely on geometric alignment, we further incorporate semantic similarity estimated during denoising, enabling more reliable correspondence modeling and robust multi-view editing. In addition, we design a selective editing pipeline that allows users to choose preferred results from multiple candidates, offering greater flexibility and user control. Extensive experiments show that CoreEditor produces high-quality, 3D-consistent edits with sharper details, significantly outperforming prior methods.
@@
3D Editing, Gaussian Splatting, Diffusion
@@ I Introduction
Figure 1: Key features of our method and visual comparison with the state-of-the-art GaussCtrl [1]. (a) Multi-view training image editing results. CoreEditor integrates geometric and semantic correspondences into the T2I diffusion model, achieving superior 3D-consistent edits compared with prior methods. (b) Rendered edited results. Leveraging consistent multi-view inputs, CoreEditor produces outputs with sharper textures and improved visual fidelity.
Recent advances in neural 3D representations, exemplified by NeRF [2] and Gaussian Splatting [3], have pushed the boundaries of photorealistic novel view synthesis, setting new benchmarks for fidelity and efficiency. Although these methods excel at reconstructing high-quality 3D scenes from multi-view images, once constructed, modifying these scenes to align with user preferences remains a challenging task. Developing such a 3D editing tool has become a critical research focus.
With the recent advances of diffusion-based text-to-image (T2I) models [4, 5, 6], most text-driven 3D editing approaches adapt these models to edit multi-view images in a zero-shot manner. However, the stochastic sampling process of diffusion models often induces cross-view inconsistencies, undermining their reliability for 3D scene optimization. The pioneering InstructNeRF2NeRF [7] addresses this issue via an iterative dataset update scheme, but suffers from slow convergence. Subsequent works [8, 1, 9] accelerate the process and enhance editing quality by jointly editing multiple views and enforcing inter-view alignment through strategies such as depth-conditioned ControlNet [6], cross-frame attention [10], and cross-view feature interpolation [9]. However, these strategies lack explicit constraints on multi-view information exchange, thereby compromising the consistency of image details. This limitation becomes particularly pronounced under significant viewpoint variations and complex occlusions, often leading to inadequate edits and degraded details. As shown in Fig. 1, GaussCtrl [1] produces inconsistent and low-quality multi-view edits, resulting in rendering results with blurry texture.
@@
To address the first one, we design a geometry and semantic co-supported approach to extract the multi-view correspondences. The key insight here is that semantically similar patches can also be involved in attention to improve consistency. As shown in Fig. 1 (a), although the left eye is occluded in the rightmost image, the accessible right eye is also expected to be visually consistent with the left one. Therefore, we are inspired to enrich the sparse geometric correspondences with semantic information. Specifically, in regions where geometric correspondences are unavailable, additional correspondences are calculated based on the diffusion feature similarity. With this comprehensive correspondence, CCA generates more plausible multi-view edits. Regarding the second problem, we introduce a selective editing pipeline, where users are allowed to select their preferred editing pattern from the per-view editing results. The selected edit is then injected into the diffusion model by a Reference Attention (RA). This approach ensures preliminary alignment of the global editing style, enabling CCA to focus solely on local consistency. Moreover, CoreEditor can generate diverse yet faithful 3D edits by selecting different per-view editing patterns.
@@
  • We propose a geometry and semantic co-supported approach to build the multi-view correspondences, which significantly improves editing quality in complex scenes.
@@ II-A Text-driven 3D Editing
Early methods [16, 17, 18] primarily leverage vision-language models [19] for text-driven 3D stylization. However, their capabilities are often limited to modifying only the global style of the scene. Building upon the success of diffusion models, DreamFusion [20] introduces a score distillation sampling (SDS) loss for 3D generation from arbitrary text, which implicitly transfers prior knowledge from a pre-trained T2I model. The SDS loss has since been applied to 3D editing and personalization in several subsequent works [21, 22, 23, 24, 25].
InstructNerf2Nerf [7] is the first method to explicitly utilize a T2I model for this task. It addresses the issue of multi-view inconsistency by iteratively alternating between editing the training images and optimizing the 3D scene, which, however, results in a slow editing process. Following-up works improve the editing performance and speed by leveraging the explicit properties of 3D Gaussian splitting [26, 27, 28, 29], latent space optimization [30, 31, 32], personalized editing [23, 33, 34], progressive editing [35], and 3D-aware fine-tuning of the diffusion model [36, 37, 38, 39]. Depth images are frequently used in recent methods to link different views. For example, VICA-NeRF [40] and DATENeRF [8] project edited images to other views using depth information. Among them, a prevalent approach involves propagating information between different views during multi-view joint editing. GaussCtrl [1] utilizes depth as guidance for ControlNet [6] and aligns the latent code of different views. DGE [9] applies cross-frame attention to edit key views, blending the edited features based on epipolar constraints. EditSplat [29] proposes a multi-view classifier-free guidance strategy to guide the diffusion model. Despite these advances, existing strategies often lack accurate constraints during the multi-view joint editing. Thereby, they typically fail to maintain precise 3D consistency. In contrast, we propose a novel correspondence-constrained attention mechanism, where only the corresponding tokens across views can communicate with each other. This enables a reliable information exchange across views, minimizing the introduction of irrelevant content and ensuring precise 3D consistency.
@@ III-B Overview: Selective Editing Pipeline
(2) In the second step, we incorporate multi-view constraints into the diffusion process and jointly edit images in $\mathcal{I}$ to an image set $\mathcal{I}^{e}$ with consistent local details. In detail, $\mathcal{I}$ and $\mathcal{D}$ are used to build geometry and semantic co-supported correspondence (Sec. III-D). Those correspondences are integrated by introducing a Correspondence-constrained Attention (CCA) module (Sec. III-E) in the U-Net. With the modified diffusion model, we use the inversion-based editing method to get $\mathcal{I}^{e}$, which is then used to optimize $\mathcal{G}$.
@@ III-C Reference Attention (RA)
Since edited results for the same prompt can be completely different across views, posing difficulties in producing high-quality edits while keeping consistency between them, we allow the users to select their preferred editing pattern, which guides the entire editing process through attention feature injection. Specifically, we modify the SA module within the diffusion backward process, transforming it into an RA module. As presented in Fig. 3, compared with SA, the diffusion feature $F^{r}$ of $I^{r}$ serves as an additional key and value in RA, thus facilitating the alignment of the editing style. Given the multi-view input features $\mathcal{F}=\{F_{i}\}_{i=1}^{N}$ in the reference attention module, the output feature ZiZ{i} for the $i$-th view is computed as:
@@
III-D Geometry and Semantic Co-supported Correspondence
@@
Since valuable information exists only in some specific target views, we only use semantic correspondences with the highest cosine similarity exceeding a threshold value $\beta$. Other semantic correspondences are retained masked. Here, to avoid additional hyper-parameter tuning, we typically set $\beta$ to 0.9, which performs well across various scenes. Finally, a comprehensive correspondence set $\mathcal{C}$ has been established, accompanied by a mask $M$ indicating the validity of each correspondence.
@@ III-E Correspondence-constrained Attention (CCA)
We then incorporate $\mathcal{C}$ into the diffusion U-Net through the proposed CCA module. Detailedly, in both the DDIM inversion and the denoising backward process, a CCA is placed after each self-attention/reference attention module. Given the multi-view features $\mathcal{Z}=\{Z_{i}\}^{N}_{i=1}$ output by self/reference attention modules, the output token $P^{{}^{\prime}}$ from CCA can be calculated as:
@@ IV-A Experimental Setup
TABLE I: Quantitative comparison with recent methods [27, 9, 1, 29] ($_{sim}: CLIP$ similarity scores, $_{dir}: CLIP$ directional similarity scores, Met3R values [63], and User study voting rates).
@@
GaussianEditor [27] 0.244 0.086 - 7.0% 8.8%
DGE [9] 0.259 0.123 0.390 14.6% 15.0%
GaussCtrl [1] 0.257 0.128 0.372 16.2% 14.8%
EditSplat [29] 0.261 0.130 0.336 17.0% 19.4%
Ours 0.270 0.145 0.281 45.2% 42.0%
@@ IV-B Qualitative Results
Based on the results, we can conclude that the inconsistency issues of other methods are mainly reflected in two aspects: (1) Incomplete Editing: When the edited training images are highly inconsistent, the resulting 3d scenes often exhibit insufficient visual change. For instance, GaussianEditor and EditSplat fail to modify the original GS model in the “decayed skeleton horse” cases in Fig. 5. GaussCtrl exhibits similar limitations, as evidenced by its inability to successfully transform the human subject into an ape-man in Fig. 4. (2) Degraded Rendering Quality: In some scenarios, although existing methods can partially achieve the target edits, they still produce locally inconsistent multi-view edits, resulting in Blurry renderings with noticeable artifacts (highlighted by yellow dashed boxes in Fig. 4 and Fig. 5). Specifically, for the stylization of $360^{\circ}$ scenes, such as the “snow” and “Ukiyo-e” cases in Fig. 4, the inconsistency introduces foggy artifacts that significantly degrade visual quality Similarly, in the “panda” case in Fig. 5, all the competitors produce vague panda faces. In contrast, CoreEditor, equipped with the proposed CAA module and selective editing pipeline, effectively aligns multi-view edits at both global and local levels, significantly outperforming existing methods. Besides, we also present a qualitative comparison with GaussCtrl [1] using a free-viewpoint rendering video in the supplemental material. The comparison highlights that CoreEditor produces more faithful edits while substantially mitigating flickering artifacts. This further underscores the superior consistency brought by our method.
@@ IV-C Quantitative Results
The quantitative comparison with baseline methods is summarized in Tab. I. Following the previous practice [7, 1, 9], we evaluate the performance using two CLIP-based metrics computed on rendered images: the CLIP similarity score and the CLIP directional similarity score [7]. The CLIP similarity score measures the degree of alignment between the edited images and the target text prompt, while the CLIP directional similarity score assesses how well the visual changes correspond to the semantic changes implied by the text. As shown in Tab. I, CoreEditor consistently outperforms all baseline methods across both metrics, highlighting its superior ability to generate edits that are semantically faithful to the text prompt. To quantitatively evaluate the 3D consistency of CoreEditor, we employ the recently proposed Met3R metric [63], which measures feature similarity between view-warped DINO [64] features. We compute the Met3R values on the edited multi-view training images generated by DGE, GaussCtrl, EditSplat and our method. GaussianEditor is excluded from this comparison as it does not support joint multi-view image editing. The results demonstrate that CoreEditor significantly improves 3D consistency compared to baseline methods through the proposed techniques. Moreover, given the inherently subjective nature of 3D editing, we validate our method through a user study involving 50 participants and 10 editing prompts. Given the source scene, editing prompts, and the rendered videos of each method, participants were asked to select the best method based on two criteria, respectively: (1) Overall visual quality and (2) 3D consistency (mainly based on the frequency of flickering artifacts in rendered videos). The user voting rates in Tab. I indicate a clear preference for CoreEditor’s outputs, further corroborating the advantages of our approach from a human perceptual perspective.
Efficiency Analysis. Leveraging its zero-shot diffusion design, CoreEditor accomplishes multi-view joint editing with only 18 GB of GPU memory under default settings. It edits scene in about 8 minutes—faster than EditSplat (12 min) and GaussianEditor (25 min), comparable to GaussCtrl (10 min), and only slightly slower than the fastest baseline, DGE (5 min). Crucially, CoreEditor achieves these results while delivering substantial and consistent gains in editing quality, offering a favorable balance between accuracy and efficiency.
Figure 6: Results with different per-view edits $I^{r}$ selected.

### v2

# CoreEditor: Correspondence-constrained Diffusion for Consistent 3D Editing

@@ Abstract
Text-driven 3D editing is an emerging task that focuses on modifying scenes based on text prompts. Current methods often adapt pre-trained 2D image editors to multi-view observations, using specific strategies to combine information across views. However, these approaches still struggle with ensuring consistency across views, as they lack precise control over the sharing of information, resulting in edits with insufficient visual changes and blurry details. In this paper, we propose CoreEditor, a novel framework for consistent text-to-3D editing. At the core of our approach is a novel correspondence-constrained attention mechanism, which enforces structured interactions between corresponding pixels that are expected to remain visually consistent during the diffusion denoising process. Unlike conventional wisdom that relies solely on scene geometry, we enhance the correspondence by incorporating semantic similarity derived from the diffusion denoising process. This combined support from both geometry and semantics ensures a robust multi-view editing process. Additionally, we introduce a selective editing pipeline that enables users to choose their preferred edits from multiple candidates, creating a more flexible and user-centered 3D editing process. Extensive experiments demonstrate the effectiveness of CoreEditor, showing its ability to generate high-quality 3D edits, significantly outperforming existing methods. Our code is available at https://github.com/czvvd/CoreEditor.
@@
3D Editing, Gaussian Splatting, Diffusion.
@@ I Introduction
Figure 1: Key features of our method and visual comparison with the recent GaussCtrl [61] method. (a) Visual comparison of edited multi-view training images. CoreEditor integrates geometric and semantic correspondences into the T2I diffusion model, ensuring 3D-consistent edits. (b) Visual comparison of rendered edited results. With consistent multi-view images, CoreEditor generates results with sharper textures.
Recent years have witnessed remarkable progress in neural 3D representations, with pioneering works like NeRF [35] and Gaussian Splatting [23] achieving photorealistic novel view synthesis. These methods excel at reconstructing high-quality 3D scenes from multi-view images. However, once constructed, modifying these scenes to align with user preferences remains a challenging task. Developing such a 3D editing tool has become a critical research focus.
With recent successes of diffusion-based text-to-image (T2I) models [19, 43, 65], most text-driven 3D editing methods typically leverage these T2I models to edit multi-view images in a zero-shot manner. Yet, the stochastic nature of diffusion models often leads to inconsistent editing results across views, producing imprecise 3D editing with blurry textures. To overcome the inconsistency issue, the pioneering InstructNerf2Nerf [14] proposes an iterative dataset update scheme but suffers from slow optimization. In an effort to accelerate the editing process, some recent works [41, 61, 9] jointly edit multiple views and align the editing pattern between these views through various strategies, such as depth-conditioned ControlNet [65], cross-frame attention [60], and cross-view feature interpolation [9]. However, these strategies lack precise constraints on the direction of multi-view information exchange, thereby compromising the consistency of local image details, especially under significant viewpoint variations and complex 3D scene occlusions. As presented in Fig. 1 (a), GaussCtrl [61] produces inconsistent and low-quality multi-view edits, resulting in results with blurry texture.
@@
To address the first one, we design a geometric and semantic co-supported approach to extract the multi-view correspondences. The key insight here is that semantically similar patches can also be involved in attention to improve consistency. As shown in Fig. 1 (a), although the left eye is occluded in the rightmost image, the accessible right eye is also expected to be visually consistent with the left one. Therefore, we are inspired to enrich the sparse geometric correspondences with semantic information. Specifically, in regions where geometric correspondences are unavailable, additional correspondences are calculated based on the diffusion feature similarity. With this comprehensive correspondence, CCA generates more plausible multi-view edits. Regarding the second problem, we introduce a selective editing pipeline, where users are allowed to select their preferred editing pattern from the per-view editing results. The selected edit is then injected into the diffusion model by a Reference Attention (RA). This approach ensures preliminary alignment of the global editing style, enabling CCA to focus solely on local consistency. Moreover, CoreEditor can generate diverse yet faithful 3D edits by selecting different per-view editing patterns.
@@
  • We propose a geometric and semantic co-supported approach to build the multi-view correspondences, which significantly improves editing quality in complex scenes.
@@ II-A Text-driven 3D Editing
Early methods [34, 54, 55] primarily leverage vision-language models [40] for text-driven 3D stylization. However, their capabilities are often limited to modifying only the global style of the scene. Building upon the success of diffusion models, DreamFusion [39] introduces a score distillation sampling (SDS) loss for 3D generation from arbitrary text, which implicitly transfers prior knowledge from a pre-trained T2I model. The SDS loss has since been applied to 3D generation and editing in several subsequent works [45, 69, 68, 36, 6, 17, 27].
InstructNerf2Nerf [14] is the first method to explicitly utilize a T2I model for this task. It addresses the issue of multi-view inconsistency by iteratively alternating between editing the training images and optimizing the 3D scene, which, however, results in a slow editing process. Following-up works improve the editing performance and speed by leveraging the explicit properties of 3D Gaussian splitting [56, 11, 58, 29], latent space optimization [10, 16, 24], personalized editing [68, 47, 15], progressive editing [8], and 3D-aware fine-tuning of the diffusion model [33, 4, 7, 62]. Depth images are frequently used in recent methods to link different views. For example, VICA-NeRF [13] and DATENeRF [41] project edited images to other views using depth information. Among them, a prevalent approach involves propagating information between different views during multi-view joint editing. GaussCtrl [61] utilizes depth as guidance for ControlNet [65] and aligns the latent code of different views. DGE [9] applies cross-frame attention to edit key views, blending the edited features based on epipolar constraints. EditSplat [29] proposes a multi-view classifier-free guidance strategy to guide the diffusion model. InterGSEdit [59] enforces multi-view consistency through a dynamic gating mechanism in cross-view attention; however, it still relies on a set of edited key frames, which can be inconsistent. Despite these advances, existing strategies often lack accurate constraints during the multi-view joint editing. Thereby, they typically fail to maintain precise 3D consistency. In contrast, we propose a novel correspondence-constrained attention mechanism, where only the corresponding tokens across views can communicate with each other. This enables a reliable information exchange across views, minimizing the introduction of irrelevant content and ensuring precise 3D consistency.
@@ III-B Overview: Selective Editing Pipeline
(2) In the second step, we incorporate multi-view constraints into the diffusion process and jointly edit images in $\mathcal{I}$ to an image set $\mathcal{I}^{e}$ with consistent local details. In detail, $\mathcal{I}$ and $\mathcal{D}$ are used to build geometric and semantic co-supported correspondence (Sec. III-D). Those correspondences are integrated by introducing a Correspondence-constrained Attention (CCA) module (Sec. III-E) in the U-Net. With the modified diffusion model, we use the inversion-based editing method to get $\mathcal{I}^{e}$, which is then used to optimize $\mathcal{G}$.
@@ III-C Reference Attention (RA)
Since edited results for the same prompt can be completely different across views, posing difficulties in producing high-quality edits while keeping consistency between them, we allow the users to select their preferred editing pattern, which guides the entire editing process through attention feature injection. Specifically, we modify the SA module within the diffusion backward process, transforming it into an RA module. As presented in Fig. 3, compared with SA, the diffusion feature $F^{r}$ of $I^{r}$ serves as an additional key and value in RA, thus facilitating the alignment of the editing style. Given the multi-view input features $\mathcal{F}=\{F_{i}\}_{i=1}^{N}$ in the reference attention module, the output feature ZiZ{i} for the $i$-th view is computed as follows:
@@
III-D Geometric and Semantic Co-supported Correspondence
@@
Since valuable information exists only in some specific target views, we only use semantic correspondences with the highest cosine similarity exceeding a threshold value $\beta$. Other semantic correspondences remain masked. Here, to avoid additional hyper-parameter tuning, we typically set $\beta$ to 0.9, which performs well across various scenes. Finally, a comprehensive correspondence set $\mathcal{C}$ has been established, accompanied by a mask $M$ indicating the validity of each correspondence.
@@ III-E Correspondence-constrained Attention (CCA)
We then incorporate $\mathcal{C}$ into the diffusion U-Net through the proposed CCA module. Specifically, in both the DDIM inversion and the denoising backward process, a CCA is placed after each self-attention/reference attention module. Given the multi-view features $\mathcal{Z}=\{Z_{i}\}^{N}_{i=1}$ output by self/reference attention modules, the output token $P^{{}^{\prime}}$ from CCA can be calculated as follows:
@@ IV-A Experimental Setup
TABLE I: Quantitative comparison with recent methods [11, 9, 61, 29] ($_{sim}: CLIP$ similarity scores, $_{dir}: CLIP$ directional similarity scores, Met3R values [1], and User study voting rates). For CLIP-based metrics, we report both the mean and median values across all evaluation cases (shown as mean / median).
@@
GaussianEditor [11] 0.244 / 0.235 0.086 / 0.081 - 7.0% 8.8%
DGE [9] 0.259 / 0.242 0.123 / 0.137 0.390 14.6% 15.0%
GaussCtrl [61] 0.257 / 0.251 0.128 / 0.126 0.372 16.2% 14.8%
EditSplat [29] 0.261 / 0.252 0.130 / 0.132 0.336 17.0% 19.4%
Ours 0.270 / 0.259 0.145 / 0.141 0.281 45.2% 42.0%
@@ IV-B Qualitative Results
Based on the results, we can conclude that the inconsistency issues of other methods are mainly reflected in two aspects: (1) Incomplete Editing: When the edited training images are highly inconsistent, the resulting 3d scenes often exhibit insufficient visual change. For instance, GaussianEditor and EditSplat fail to modify the original GS model in the “decayed skeleton horse” cases in Fig. 4. GaussCtrl exhibits similar limitations, as evidenced by its inability to successfully transform the human subject into an ape-man in Fig. 5. (2) Degraded Rendering Quality: In some scenarios, although existing methods can partially achieve the target edits, they still produce locally inconsistent multi-view edits, resulting in blurry renderings with noticeable artifacts (highlighted by yellow dashed boxes in Fig. 5 and Fig. 4). Specifically, for the stylization of $360^{\circ}$ scenes, such as the “snow” and “Ukiyo-e” cases in Fig. 5, the inconsistency introduces foggy artifacts that significantly degrade visual quality. Similarly, in the “panda” case in Fig. 4, all the competitors produce vague panda faces. In contrast, CoreEditor, equipped with the proposed CAA module and selective editing pipeline, effectively aligns multi-view edits at both global and local levels, significantly outperforming existing methods. Besides, we also present a qualitative comparison with GaussCtrl [61] using a free-viewpoint rendering video in the supplemental material. The comparison highlights that CoreEditor produces more faithful edits while substantially mitigating flickering artifacts. This further underscores the superior consistency brought by our method.
Figure 6: Results with different per-view edits $I^{r}$ selected.
Figure 9: Visual comparison with PDS [27] and Morpheus [62]. We provide results rendered from two views for each edited scene. Blurry regions are highlighted with yellow dash boxes.
Figure 10: Multi-view editing results produced by InstructPix2Pix [3] and its variant augmented with the proposed CoreEditor, which improves multi-view consistency.
Figure 11: Qualitative comparison with recent methods [9, 61, 29] on prompts requiring geometric changes.
Figure 12: Effect of each component on editing cases that require geometric changes. We show the rendered RGB and depth images for each view.
@@ IV-C Quantitative Results
The quantitative comparison with baseline methods is summarized in Tab. I. Following the previous practice [14, 61, 9], we evaluate the performance using two CLIP-based metrics computed on rendered images: the CLIP similarity score and the CLIP directional similarity score [14]. For each scene and prompt, CLIP scores are first averaged across rendered views to obtain a single score. We then report both the mean and the median across all evaluation cases, as shown in Tab. I. The CLIP similarity score measures the degree of alignment between the edited images and the target text prompt, while the CLIP directional similarity score assesses how well the visual changes correspond to the semantic changes implied by the text. As shown in Tab. I, CoreEditor consistently outperforms all baseline methods across both metrics, highlighting its superior ability to generate edits that are semantically faithful to the text prompt. To quantitatively evaluate the 3D consistency of CoreEditor, we employ the recently proposed Met3R metric [1], which measures feature similarity between view-warped DINO [38] features. We compute the Met3R values on the edited multi-view training images generated by DGE, GaussCtrl, EditSplat and our method. GaussianEditor is excluded from this comparison as it does not support joint multi-view image editing. The results demonstrate that CoreEditor significantly improves 3D consistency compared to baseline methods through the proposed techniques. Moreover, given the inherently subjective nature of 3D editing, we validate our method through a user study involving 50 participants and 10 editing prompts. Given the source scene, editing prompts, and the rendered videos of each method, participants were asked to select the best method based on two criteria, respectively: (1) Overall visual quality and (2) 3D consistency (mainly based on the frequency of flickering artifacts in rendered videos). The user voting rates in Tab. I indicate a clear preference for CoreEditor’s outputs, further corroborating the advantages of our approach from a human perceptual perspective.
Efficiency Comparison. Since the diffusion model in CoreEditor operates in a zero-shot manner, our multi-view joint editing only requires about 18 GB of GPU memory under the default settings. In terms of runtime, CoreEditor typically completes 3D scene editing within 8 minutes. While DGE achieves faster processing (5 minutes), other methods are slower than CoreEditor: GaussCtrl (10 minutes), EditSplat (12 minutes), and GaussianEditor (25 minutes). Despite not being the fastest, CoreEditor delivers significant improvements in editing quality while maintaining a computationally efficient design, avoiding additional overhead.